Calculate the mass of nitrogen present in a volume of 3000 cm3 if the temperature of the gas is 22.0 ∘C and the absolute pressure 2.00×10−13 atm is a partial vacuum easily obtained in laboratories. The molar mass of nitrogen (N2) is 28.0 g/mol . Express your answer in kg
Answer #1

Given:
P = 2.0*10^-13 atm
V = 3000 cm^3
= 3000 mL
= (3000/1000) L
= 3 L
T = 22.0 oC
= (22.0+273) K
= 295 K

find number of moles using:
P * V = n*R*T
0 atm * 3 L = n * 0.08206 atm.L/mol.K * 295 K
n = 2.479*10^-14 mol

Molar mass of N2 = 28.02 g/mol

use:
mass of N2,
m = number of mol * molar mass
= 2.479*10^-14 mol * 28.02 g/mol
= 6.945*10^-13 g
= 6.945*10^-16 Kg
Answer: 6.94*10^-16 Kg